OpenStudy (anonymous):

Write a rule for the nth term of the sequence 37, 32, 27, 22, ...

OpenStudy (anonymous):

37-5n

OpenStudy (anonymous):

It would be tn=42 - 5*n where tn is the nth term.

OpenStudy (anonymous):

A. a n = -5n + 42 B. a n = -3n + 40 C. a n = 5n + 32 D. a n = 5n - 42

OpenStudy (anonymous):

so A. n = -5n + 42

OpenStudy (anonymous):

5n - 42? xD

OpenStudy (anonymous):

ah. thank you!

OpenStudy (anonymous):

42 - 5n can also be written as -5n + 42

OpenStudy (anonymous):

gotcha, thanks :)

OpenStudy (anonymous):

Mathematica agrees with Harkirat.\[\text{FindSequenceFunction}[\{37,32,27,22\},n]=42-5 n \]\[\text{Table}[42-5 n,\{n,8\}]=\{37,32,27,22,17,12,7,2\} \]

OpenStudy (anonymous):

Thank you for the medal.

OpenStudy (anonymous):

You got it!